# A Minimalist Optimizer Design for LLM Pretraining

### v2

@@ 2.2 Momentum in the last layer
First-order momentum ($m^{t}$ in (2.3)) has been shown to effectively reduce the variance toward the true gradients; see liu2020improved; cutkosky2020momentum. Although second-order momentum is used in adaptive optimizers such as Adam and RMSprop, more recent optimizers such as Muon have demonstrated remarkable success without second-order momentum. It is worth noticing that momentum is the major factor that introduces memory overhead in optimizer states, therefore we only consider first-order momentum in our optimizer design. Based on a minimalist optimizer design principle, we plan to investigate whether first-order momentum can be selectively used among different layers, while still leading to performance improvements without significantly increasing memory overhead.
Layer-wise gradient variance. Existing works already show the effectiveness of momentum in reducing variances (liu2020improved; cutkosky2020momentum). Naturally, if the gradients for different layers have different variances, we could achieve both memory efficiency and variance reduction by applying momentum to the layers with large variances. In the next paragraphs, we identify the layer-wise gradient variances and determine the most important layer that provides the largest performance gain when incorporating momentum; we also justify the effectiveness of different momentum parameters for different layers using a convergence analysis theory in Theorem 2.1.
We first perform a simple experiment on LLaMA 130M to check the gradient variance of different layers. To estimate the gradient variance, one needs to obtain the full gradient (by using the entire training dataset) which is not feasible in practice. Instead, we take a much larger training data batch as input [^4] For example, we take the training batch size as $32$ and the large batch size as $512$. to estimate the true gradient. The experiment results of running SGD with column-wise normalization (SGD-col-norm) and SGD-col-norm with last layer momentum are shown in Figure 4. Interestingly, we observe that the variance of the last layer is the largest among all the layers during the training time, necessitating a specific treatment for the last layer to reduce the variance.
@@
Figure 4: Estimated variance of the stochastic gradients (and momentum when applicable) for different layers in two methods (smoothed by 50 iterations window). We observe that when running SGD with column-wise normalization (SGD-col-norm, left plot), the variance of the last layer (lm_head) is largest for most of the time, following by the variance of the first layer (embedding) and other layers. After applying momentum to the last layer (SGD-col-norm-mmt-last, right plot), the variance of the momentum of last layer (lm_head momentum) decreases to a very low level. Interestingly, the variance of the first layer in plot (b) is also smaller than the one in plot (a).
Next, we show theoretically that momentum helps the most for the layers with larger gradient variances. We inspect the theoretical property of applying SGD with momentum (SGD-M) to the LLM optimization problem (2.1). Note that here we did not consider column-wise normalization for ease of theoretical analysis. Consider the following SGD-M algorithm to solve (2.1) ($m_{l}^{0}=0$ and $\theta^{1}$ is randomly initialized):
$m_{l}^{t}=\beta_{l}m_{l}^{t-1}+(1-\beta_{l})g^{t},\ g^{t}=\nabla_{\theta_{l}}\ell(\theta^{t};\xi_{t})$ (2.7)
@@
where t=1,2,,Tt=1,2,...,T is the iteration counter; l=1,,Ll=1,...,L represents different layers, i.e. we assume different layers contain different momentum with different hyperparameters. We have the following theoretical result (see Appendix A.10 for the proof).
@@
Suppose $\ell(\theta)$ in (2.1) is lower bounded by $\ell^{*}$, $\gamma$-smooth (i.e. $\nabla\ell(\theta)$ is Lipschitz continuous with constant $\gamma$), also the stochastic gradient is unbiased $\mathbb{E}_{\xi_{t}}\nabla_{\theta_{l}}\ell(\theta^{t};\xi_{t})=\nabla_{\theta_{l}}\ell(\theta^{t})$ and with bounded variance: $\mathbb{E}_{\xi_{t}}\|\nabla_{\theta_{l}}\ell(\theta^{t};\xi_{t})-\nabla_{\theta_{l}}\ell(\theta^{t})\|^{2}\leq\sigma_{l}^{2}$ for all l=1,,Ll=1,...,L and t=0,,T1t=0,...,T-1. With appropriate choice of hyperparameters $\eta_{l}\geq\eta$ (see (A.20) and (A.10)) and $\beta_{l}\leq 1-\delta$ ($\delta$ is an absolute constant), we have the following convergence result for update (2.7):
$\frac{1}{T}\sum_{t=1}^{T}\sum_{l=1}^{L}\mathbb{E}\|\nabla_{l}^{t}\|^{2}\leq\frac{2L\gamma^{3/2}\mathbb{E}\Delta_{1}}{\delta^{2}\sqrt{T}}+\sum_{l=1}^{L}\left(\frac{1-\beta_{l}}{1+\beta_{l}}\frac{L\sqrt{\gamma}}{4\sqrt{T}}+\frac{L\gamma^{3/2}}{2\sqrt{T}}+\frac{1-\beta_{l}}{\beta_{l}^{3}}\frac{\gamma^{2}}{4LT}\right)\frac{\sigma_{l}^{2}}{\delta^{2}}$ (2.8)
@@
The left-hand side of (2.8) is the expectation of the stochastic gradient, measuring the stationarity. The right-hand side suggests an overall $\mathcal{O}(1/\sqrt{T})$ rate of convergence in terms of the iteration number $T$, meanwhile the summation indicates the error to each of the layer is aggregated, whose magnitude depends on the variance level $\sigma_{l}$ of the specific layer.
Theorem 2.1 suggests that by taking different $\beta_{l}$ for different layers l=1,,Ll=1,...,L, the convergence could be improved. The second term of the right-hand side of (2.8) takes the form
@@
where $x$ represents the layer-wise momentum hyperparameter $\beta_{l}$. It is straightforward to verify that $f(x)$ is decreasing in $(0,1-\delta]$, therefore an optimal strategy is to pick $\beta_{l}=1-\delta$. However from a memory-efficient point of view, only $\beta_{l}=0$ saves the memory of momentum.
@@
Proof Sketch for Theorem 2.1. The proof follows liu2020improved, with the major difference being that liu2020improved proved the case with one single layer ($L=1$) and we extend it to multi-layer setting ($L>1$) where each layer has difference variance level.
@@
$z_{l}^{t}:=\begin{cases}\theta_{l}^{t}&t=1\\ \frac{1}{1-\beta_{l}}\theta_{l}^{t}-\frac{\beta_{l}}{1-\beta_{l}}\theta_{l}^{t-1}&t\geq 2\end{cases}$ (2.9)
we can verify two important properties: $z_{l}^{t+1}-z_{l}^{t}=-\eta_{l}g_{l}^{t}$ and $z_{l}^{t}-\theta_{l}^{t}=-\frac{\beta_{l}}{1-\beta_{l}}\eta_{l}m_{l}^{t-1}$ (Lemma A.3). Using $\gamma$-Lipschitz smooth, we can expand the function value difference via
$\mathbb{E}_{\xi_{t}}\left[\ell\left(z^{t+1}\right)\right]\leq\ell\left(z^{t}\right)+\underbrace{\mathbb{E}_{\xi_{t}}\left[\left\langle\nabla\ell\left(z^{t}\right),z^{t+1}-z^{t}\right\rangle\right]}_{(a)}+\frac{\gamma}{2}\underbrace{\mathbb{E}_{\xi_{t}}\left[\left\|z^{t+1}-z^{t}\right\|^{2}\right]}_{(b)}.$
Expanding the term (a) on the right-hand side using the update using the above properties of $z_{l}^{t}$, we will arrive at a $\mathcal{O}(\sum_{t=1}^{T}\eta_{t}\sum_{l=1}^{L}\mathbb{E}\|\nabla_{l}^{t}\|^{2})$ term (which is the left-hand side of (2.8)), and some addition term. The additional term, combining with (b), will be bounded via Lemma A.1 and A.2, and together form the right hand side of (2.8).
@@ 3 The SCALE Algorithm
The above investigation leads to the following simple optimizer, SCALE, detailed in Algorithm 1. The design of SCALE is motivated by empirical insights from our preceding experiments, which highlighted the importance of stabilizing updates in the last layer and controlling gradient scale via column-wise normalization. Accordingly, SCALE integrates two components: column-wise normalization of gradients and momentum updates restricted to the last layer. Despite its simplicity, SCALE is highly effective and requires only minimal modifications to standard implementations of Adam—typically just a few additional lines of code. As we will demonstrate, this lightweight design allows SCALE to outperform state-of-the-art baselines such as Adam and achieve performance competitive with Stable-SPAM and Muon, while using substantially less memory.
@@
   Compute the stochastic gradient $g_{l}^{t}:=\frac{1}{B}\sum_{b=1}^{B}\nabla_{\theta_{l}}\ell(\theta^{t};\xi_{t,b})$;
@@
Connection to existing works. We now give a discussion on the connection of the proposed method to existing works. The proposed algorithm utilizes column-wise normalization, which is also discussed in pethick2025scion. pethick2025scion uses momentum for all the layers and applies column-wise normalization as an alternative to singular-value normalization only for the last layer. Our method also differs from SWAN (ma2024swan) in the following aspects: first, SWAN utilizes both the row-wise and singular-value normalization while our method only utilizes column-wise normalization, indicating certain redundancy in existing works that apply multiple normalizations; second, SWAN applies Adam for the embedding and LM head layers (the first and last layers) which increases the memory overhead, whereas our approach only introduces first-order momentum for the last layer. We summarize the techniques used in different papers in Table 4.
Convergence of Algorithm 1. We remark here that the convergence provided in pethick2025scion could be adopted to provide a convergence analysis for Algorithm 1, since Algorithm 1 differs from the general framework of pethick2025scion only by the fact that Algorithm 1 applies different momentum for different layers. However, directly applying such kinds of analysis will not be able to explain why using either normalization or momentum is advantageous as compared to vanilla SGD, we choose not to present such a proof.
@@
Table 4: A summary of building components of related methods. “Sign”, “Col-wise”, “Row-wise” and “Singular-val” correspond to four normalizations in (2.6), respectively. “1st order EMA” and “2nd order EMA” stand for first and second order EMA/momentum. The last column records the memory (GB) of weights and optimizer states required for applying the corresponding methods for LLaMA 7B training (see Appendix A.1 for the details of the estimation).
@@ 4 Experiments
In this section, we test the SCALE algorithm for LLM pretraining. We test on pretraining LLaMA (60M, 130M, 350M, 1B, 7B) models on the C4 (Colossal Clean Crawled Corpus) dataset (raffel2020exploring) and report evaluation perplexity as our metric. For all ours experiments, we follow the hyperparameter settings of zhao2024galore, see Appendix A.2 for more details.
Baselines. We compare with Adam and its stabilized version, noted as Adam (Stable-SPAM), which performs momentum resets and clips spiked gradients (huang2025stablespam). Among memory efficient optimizers, we compare with Muon, GaLore, Fira, APOLLO(-Mini) and SWAN (see Section 1.1 for a more detailed introduction of these works). GaLore, Fira and APOLLO(-Mini) compress the Adam states by projecting the gradients to low rank. SWAN is an optimizer that uses both singular-value and row-wise normalization, except for the first and last layers [^5] We copy SWAN’s result from ma2024swan since we cannot replicate them due to code unavailability..
@@
Results for 60M–1B models. We report the results of the evaluation perplexity in Table 5. We can see that, despite its simplicity, SCALE outperforms existing memory efficient optimizers, also (nearly) matches the performance of Adam (Stable-SPAM) and Muon, especially for larger models, with only 35–65% of the memory. We also contrast the performance with the memory consumed in Figure 1, where we can see that the proposed method is indeed at the Pareto frontier for optimal memory usage while maintaining the (near) state-of-the-art performance. This makes it a strong candidate for large-scale pretraining under memory constraints [^6] Notice that we construct our optimizer step by step via a minimalist design, and we already conduct the ablations throughout our journey (see Table 2 and 3, also Figure 4). We do not conduct further ablation studies in this section, but we include experiments on throughput analysis, additional architectures, overtraining, finetuning and learning rate sensitivity in the Appendix sections A.3, A.4, A.5, A.6 and A.7, respectively..
@@
Table 5: Experiment results for pretraining LLaMA models on C4 dataset. The result marked is from ma2024swan. The results marked for model sizes 60M-350M are from glentis2025scalable. All models are trained up-to the Chinchilla optimal number of tokens (hoffmann2022training). For Fira and APOLLO 1B runs we encountered training divergence with their default learning rates, results reported are with reduced learning rate. Among the memory-efficient optimizers we highlight the best-performing for each model size in terms of perplexity and memory. Note that SWAN does not provide a 1B model result with Chinchilla optimal tokens, but only for 13B tokens. In that setting we also achieve a superior perplexity of 14.25 (2.81G memory), compared to 14.42 (3.20G memory) of SWAN.
@@
Table 6: Results for pretraining 7B LLaMA model on C4 dataset. From zhu2025apollo.

### v3

@@ 2.2 Momentum in the last layer
First-order momentum ($m^{t}$ in (3)) has been shown to effectively reduce the variance toward the true gradients; see Liu et al. (2020); Cutkosky and Mehta (2020). Although second-order momentum is used in adaptive optimizers such as Adam and RMSprop, more recent optimizers such as Muon have demonstrated remarkable success without second-order momentum. It is worth noticing that momentum is the major factor that introduces memory overhead in optimizer states, therefore we only consider first-order momentum in our optimizer design. Based on a minimalist optimizer design principle, we plan to investigate whether first-order momentum can be selectively used among different layers, while still leading to performance improvements without significantly increasing memory overhead.
Layer-wise gradient variance. Existing works already show the effectiveness of momentum in reducing variances (Liu et al., 2020; Cutkosky and Mehta, 2020). Naturally, if the gradients for different layers have different variances, we could achieve both memory efficiency and variance reduction by applying momentum to the layers with large variances. In the next paragraphs, we identify the layer-wise gradient variances and determine the most important layer that provides the largest performance gain when incorporating momentum; we also justify the effectiveness of different momentum parameters for different layers using a convergence analysis theory in Theorem 2.1.
We first perform a simple experiment on LLaMA 130M to check the gradient variance of different layers. To estimate the gradient variance, one needs to obtain the full gradient (by using the entire training dataset) which is not feasible in practice. Instead, we take a much larger training data batch as input [^4] For example, we could take the small training batch size as $32$ and the large batch size as $512$. to estimate the true gradient. The experiment results of running SGD with column-wise normalization (SGD-col-norm) and SGD-col-norm with last layer momentum are shown in Figure 4. Interestingly, we observe that the variance of the last layer is the largest among all the layers during training, necessitating a specific treatment for the last layer to reduce the variance.
@@
Figure 4: Estimated variance of the stochastic gradients (and momentum when applicable) for different layers in two methods (smoothed by 50 iterations window). We observe that when running SGD with column-wise normalization (SGD-col-norm, left plot), the variance of the last layer (lm_head) is largest for most of the time, followed by the variance of the first layer (embedding) and other layers. After applying momentum to the last layer (SGD-col-norm-mmt-last, right plot), the variance of the momentum of last layer (lm_head momentum) decreases to a very low level. Interestingly, the variance of the first layer in plot (b) is also smaller than the one in plot (a).
Next, we show theoretically that momentum helps the most for the layers with larger gradient variances. We inspect the theoretical property of applying SGD with momentum (SGD-M) to the LLM optimization problem (1). Note that here we did not consider column-wise normalization for ease of theoretical analysis. Consider the following SGD-M algorithm to solve (1) ($m_{l}^{0}=0$ and $\theta^{1}$ is randomly initialized):
$m_{l}^{t}=\beta_{l}m_{l}^{t-1}+(1-\beta_{l})g^{t},\ g^{t}=\nabla_{\theta_{l}}\ell(\theta^{t};\xi_{t})$ (7)
@@
where t=1,2,,Tt=1,2,...,T is the iteration counter; l=1,,Ll=1,...,L represents different layers, i.e. we assume different layers contain different momentum with different hyperparameters. We have the following theoretical result (see Appendix N for the proof).
@@
Suppose $\ell(\theta)$ in (1) is lower bounded by $\ell^{*}$, $\gamma$-smooth (i.e. $\nabla\ell(\theta)$ is Lipschitz continuous with constant $\gamma$), also the stochastic gradient is unbiased $\mathbb{E}_{\xi_{t}}\nabla_{\theta_{l}}\ell(\theta^{t};\xi_{t})=\nabla_{\theta_{l}}\ell(\theta^{t})$ and with bounded variance: $\mathbb{E}_{\xi_{t}}\|\nabla_{\theta_{l}}\ell(\theta^{t};\xi_{t})-\nabla_{\theta_{l}}\ell(\theta^{t})\|^{2}\leq\sigma_{l}^{2}$ for all l=1,,Ll=1,...,L and t=0,,T1t=0,...,T-1. With appropriate choice of hyperparameters $\eta_{l}$ (defined in (N)) and $\beta_{l}\leq 1-\delta$ ($\delta$ is an absolute constant), we have the following convergence result for update (7):
$\begin{split}&\frac{1}{T}\sum_{t=1}^{T}\sum_{l=1}^{L}\mathbb{E}\|\nabla_{l}^{t}\|^{2}\leq\frac{2L\gamma^{3/2}\mathbb{E}\Delta_{1}}{\delta^{2}\sqrt{T}}\\ &+\sum_{l=1}^{L}\left(\frac{1-\beta_{l}}{1+\beta_{l}}\frac{L\sqrt{\gamma}}{4\sqrt{T}}+\frac{L\gamma^{3/2}}{2\sqrt{T}}+\frac{1-\beta_{l}}{\beta_{l}^{3}}\frac{\gamma^{2}}{4LT}\right)\frac{\sigma_{l}^{2}}{\delta^{2}}\end{split}$ (8)
@@
The left-hand side of (8) is the expectation of the stochastic gradient, measuring the stationarity. The right-hand side suggests an overall $\mathcal{O}(1/\sqrt{T})$ rate of convergence in terms of the iteration number $T$, meanwhile the summation indicates the error to each of the layer is aggregated, whose magnitude depends on the variance level $\sigma_{l}$ of the specific layer.
Theorem 2.1 suggests that by taking different $\beta_{l}$ for different layers l=1,,Ll=1,...,L, the convergence could be improved. The second term of the right-hand side of (8) takes the form
@@
where $x$ represents the layer-wise momentum hyperparameter $\beta_{l}$. It is straightforward to verify that $f(x)$ is decreasing in $(0,1-\delta]$, therefore an optimal strategy is to pick $\beta_{l}=1-\delta$. However from a memory-efficient point of view, only $\beta_{l}=0$ saves the memory of momentums.
@@
Proof Sketch for Theorem 2.1. The proof follows Liu et al. (2020), with the major difference being that Liu et al. (2020) proved the case with one single layer ($L=1$) and we extend it to multi-layer setting ($L>1$) where each layer has different variance level.
@@
$z_{l}^{t}:=\begin{cases}\theta_{l}^{t}&t=1\\ \frac{1}{1-\beta_{l}}\theta_{l}^{t}-\frac{\beta_{l}}{1-\beta_{l}}\theta_{l}^{t-1}&t\geq 2\end{cases}$ (9)
we can verify two important properties: $z_{l}^{t+1}-z_{l}^{t}=-\eta_{l}g_{l}^{t}$ and $z_{l}^{t}-\theta_{l}^{t}=-\frac{\beta_{l}}{1-\beta_{l}}\eta_{l}m_{l}^{t-1}$ (Lemma N.3). Using $\gamma$-Lipschitz smooth, we can expand the function value difference via
$\begin{split}\mathbb{E}_{\xi_{t}}\left[\ell\left(z^{t+1}\right)\right]\leq&\ell\left(z^{t}\right)+\underbrace{\mathbb{E}_{\xi_{t}}\left[\left\langle\nabla\ell\left(z^{t}\right),z^{t+1}-z^{t}\right\rangle\right]}_{(a)}\\ &+\frac{\gamma}{2}\underbrace{\mathbb{E}_{\xi_{t}}\left[\left\|z^{t+1}-z^{t}\right\|^{2}\right]}_{(b)}.\end{split}$
Expanding the term (a) on the right-hand side using the update using the above properties of $z_{l}^{t}$, we will arrive at a $\mathcal{O}(\sum_{t=1}^{T}\eta_{t}\sum_{l=1}^{L}\mathbb{E}\|\nabla_{l}^{t}\|^{2})$ term (which is the left-hand side of (8)), and some addition term. The additional term, combining with (b), will be bounded via Lemma N.1 and N.2, and together form the right hand side of (8).
@@ 3 The SCALE Algorithm
The above investigation leads to the following simple optimizer, SCALE, detailed in Algorithm 1. The design of SCALE is motivated by empirical insights from our preceding experiments, which highlighted the importance of stabilizing updates in the last layer and controlling gradient scale via column-wise normalization. Accordingly, SCALE integrates two components: column-wise normalization of the gradients (for all layers) and momentum usage restricted to the last layer. Despite its simplicity, SCALE is highly effective and requires only minimal modifications to standard implementations of Adam—typically just a few additional lines of code. As we will demonstrate, this lightweight design allows SCALE to outperform state-of-the-art baselines such as Adam and achieve performance competitive with Stable-SPAM and Muon, while using substantially less memory.
@@
[…]
@@
Connection to existing works. We now give a discussion on the connection of the proposed method to existing works. The proposed algorithm utilizes column-wise normalization, which is also discussed in Pethick et al. (2025). Pethick et al. (2025) uses momentum for all the layers and applies column-wise normalization as an alternative to singular-value normalization only for the last layer. Our method also differs from SWAN (Ma et al., 2025) in the following aspects: first, SWAN utilizes both the row-wise and singular-value normalization while our method only utilizes column-wise normalization, indicating certain redundancy in existing works that apply multiple normalizations; second, SWAN applies Adam for the embedding and LM head layers (the first and last layers) which increases the memory overhead, whereas our approach only introduces first-order momentum for the last layer. We summarize the techniques used in different papers in Table 4.
Convergence of Algorithm 1. The convergence provided in Riabinin et al. (2025, Theorem 2) could be adopted to provide a convergence analysis for Algorithm 1, since Algorithm 1 differs from the general framework of Riabinin et al. (2025, Algorithm 1) only by the fact that we assume different gradient variance $\sigma_{i}$ for different layers. Below we provide a convergence result for a generalization of the proposed Algorithm 1:
[…]
We have the following convergence [^5] Note that to show convergence we need a different set of assumptions (which directly assume smoothness and bounded variance in matrix $1\rightarrow 2$ norm layer-wisely):
Theorem 3.1.
Suppose $\ell(\theta)$ in (1) is lower bounded by $\ell^{*}$, also each layer $l$ is $L_{l}$-smooth in matrix $1\rightarrow 2$ norm: $\|\nabla_{\theta_{l}}\ell(\theta)-\nabla_{\theta_{l}}\ell(\vartheta)\|_{2\rightarrow\infty}\leq L_{l}\|\theta-\vartheta\|_{1\rightarrow 2}$, also the stochastic gradient is unbiased $\mathbb{E}_{\xi_{t}}\nabla_{\theta_{l}}\ell(\theta^{t};\xi_{t})=\nabla_{\theta_{l}}\ell(\theta^{t})$ and with bounded variance: $\mathbb{E}_{\xi_{t}}\|\left(\nabla_{\theta_{l}}\ell(\theta^{t};\xi_{t})-\nabla_{\theta_{l}}\ell(\theta^{t})\right)_{:,j}\|_{2\rightarrow\infty}^{2}\leq\nu_{l,j}^{2}$ for all l=1,,Ll=1,...,L and t=0,,T1t=0,...,T-1 and all the columns $j$. Define $\bar{\sigma}_{l}:=\left(\sum_{j=1}^{n_{l}}\nu_{l,j}^{2}\right)^{1/2}$. If $\eta_{l}=a_{l}/\sqrt{T}$, we have the following convergence result for (10) (a generalization of Algorithm 1):
$\begin{split}&\frac{1}{T}\sum_{t=1}^{T}\sum_{l=1}^{L}a_{l}\mathbb{E}\|\nabla_{l}^{t}\|_{2\rightarrow\infty}\lesssim\frac{\Delta_{1}}{T^{1/4}}\\ &+\sum_{l=1}^{L}\left[\frac{L_{l}a_{l}^{2}}{T^{3/4}}+a_{l}\left(\frac{\beta_{l}L_{l}a_{l}}{1-\beta_{l}}\frac{1}{T^{1/2}}+\bar{\sigma}_{l}\sqrt{\frac{1-\beta_{l}}{1+\beta_{l}}}\frac{1}{T^{1/4}}\right)\right],\end{split}$ (11)
where $\nabla_{l}^{t}=\nabla_{\theta_{l}}\ell(\theta^{t};\xi_{t})$ and $\Delta_{1}=\ell(\theta^{1})$-$\ell^{*}$. Also we omitted constants in $\lesssim$ notation.
@@
Table 4: A summary of building components of related methods. “Sign”, “Col-wise”, “Row-wise” and “Singular-val” correspond to four normalizations in (6), respectively. “1st order EMA” and “2nd order EMA” stand for first and second order EMA/momentum. The last column records the memory (GB) of weights and optimizer states required for applying the corresponding methods for LLaMA 7B training (see Appendix B for the details of the estimation). Note that for the special case of vector parameters (such as LayerNorm weights) all memory-efficient methods employ the Adam optimizer with negligible impact on memory.
@@ 4 Experiments
In this section, we test the SCALE algorithm for LLM pretraining. We test on pretraining LLaMA (60M, 130M, 350M, 1B, 7B) models on the C4 (Colossal Clean Crawled Corpus) dataset (Raffel et al., 2020) and report evaluation perplexity as our metric. For all ours experiments, we follow the hyperparameter settings of Zhao et al. (2024); see Appendix C for more details. We include pretraining experiments on different architectures (GPT2-Medium, Qwen2-500M and Gemma-2B) to show the generality of the method in Appendix F.
Baselines. We compare with Adam and its stabilized version, noted as Adam (Stable-SPAM), which performs momentum resets and clips spiked gradients (Huang et al., 2025). Among memory efficient optimizers, we compare with Muon, GaLore, Fira, APOLLO(-Mini) and SWAN (see Section 1.1 for a more detailed introduction of these works). GaLore, Fira and APOLLO(-Mini) compress the Adam states by projecting the gradients to low rank. SWAN is an optimizer that uses both singular-value and row-wise normalization, except for the first and last layers [^6] We copy SWAN’s result from (Ma et al., 2025) since we cannot replicate them due to code unavailability..
@@
Results for 60M–1B models. We report the results of the evaluation perplexity in Table 5. We can see that, despite its simplicity, SCALE outperforms existing memory efficient optimizers, also (nearly) matches the performance of Adam (Stable-SPAM) and Muon, especially for larger models, with only 35–65% of the memory. We also contrast the performance with the memory consumed in Figure 1, where we can see that the proposed method is at the Pareto frontier for optimal memory usage while maintaining the (near) state-of-the-art performance. This makes it a strong candidate for large-scale pretraining under memory constraints [^7] Notice that we construct our optimizer step by step via a minimalist design, and we already conduct the ablations throughout our journey (see Table 2 and 3, also Figure 4). We do not conduct further ablation studies in this section, but we include experiments on throughput analysis, additional architectures, overtraining experiment, finetuning experiments and learning rate sensitivity in the Appendix sections D, F, H, I and K, respectively. .
@@
Table 5: Experiment results for pretraining LLaMA models on C4 dataset. The result marked is from Ma et al. (2025). The results marked for model sizes 60M-350M are from Glentis et al. (2025). All models are trained up-to the Chinchilla optimal number of tokens (Hoffmann et al., 2022). For Fira and APOLLO 1B runs we encountered training divergence with their default learning rates, results reported are with reduced learning rate. Among the memory-efficient optimizers we highlight the best-performing for each model size in terms of perplexity and memory. Note that SWAN does not provide a 1B model result with Chinchilla optimal tokens, but only for 13B tokens. In that setting we also achieve a superior perplexity of 14.25 (2.81G memory), compared to 14.42 (3.20G memory) of SWAN.
@@
Table 6: Results for pretraining 7B LLaMA model on C4 dataset. From Zhu et al. (2025).